• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 21:31:05 on October 09, 2009

    Entry number 294074

    keyword=Run plan update (21:20pm 10/09/2009)

    Run Plan (Please HALOG it if you find something strange or wrong; Call RC if the problem is not resolved within 10 minutes)

    1. 21:00-23:00 Dustin McNulty and Rupesh Silwal will perform a Q2 measurement, Aluminum dummy measurements and HRS optics data.

    2. After Dustin and Rupesh have completed, Alexandre Camsonne will be working on diagnosing the high Compton rates. This work maybe invasive to taking HAPPEx production data.

    3. Take production data! Watch the lumi widths to see if anything changes. Please check the histograms on the online data monitor with the reference ones on the shift how-to page (http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#HPXStartingonlineanalysis).

    4. Currently the Compton rates are too high! If Alexandre is successful in reducing the rates to an acceptancable level:

    5. Change the half-wave plate state when we get to 1M good pairs. Follow the instructions for doing this and running the flipper script here: http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#ChangingHalfWavePlate

    6. Run spot++ every couple of hours and log it. Change the prescale factor PS8 to 1 and PS1 and 2 to 99999.

    7. Beam Time Accounting: accept the beam unless the Compton people say that they just cannot use it. Then record it as BNA.